1. Michaelis and Menten examined how the velocity of enzyme catalyzed reactions change with substrate concentration. Which of the following is (are) common to all enzyme catalyzed reactions?

  1. Velocity is insensitive to changes in [substrate] at all substrate concentrations.
  2. Km is the [substrate] required to reach 50% of Vmax.
  3. Velocity is responsive to changes in [substrate] when the Km > [substrate].
  4. Velocity is insensitive to [substrate] when [substrate] is much greater than Km.
  5. Velocity reaches 90% of Vmax when [substrate] is 10X Km.
Answer #1

Except OPTION 'A' all other optoions are true.

MICHAELIS -MENTEN EQN AND CORRESPONDING GENERAL REACTION SCHMES ARE BELOW

When [s] >>Km, V0= vmax (statement 4)

When [S]<<Km, V0 = Vmax[S]/Km (statement 3)

When [S]=Km V0= Vmax/2 , so 50% of Vmax (statement 2)

When S= 10*Km  V0 = 10/11(Vmax) = 0.90(Vmax) , so 90% of Vmax (statement 5)